# Learning without Exact Guidance: Updating Large-scale High-resolution Land Cover Maps from Low-resolution Historical Labels

### 08 Mar 2024 v2

@@ Abstract
Large-scale high-resolution (HR) land-cover mapping is a vital task to survey the Earth’s surface and resolve many challenges facing humanity. However, it is still a non-trivial task hindered by complex ground details, various landforms, and the scarcity of accurate training labels over a wide-span geographic area. In this paper, we propose an efficient, weakly supervised framework (Paraformer) to guide large-scale HR land-cover mapping with easy-access historical land-cover data of low resolution (LR). Specifically, existing land-cover mapping approaches reveal the dominance of CNNs in preserving local ground details but still suffer from insufficient global modeling in various landforms. Therefore, we design a parallel CNN-Transformer feature extractor in Paraformer, consisting of a downsampling-free CNN branch and a Transformer branch, to jointly capture local and global contextual information. Besides, facing the spatial mismatch of training data, a pseudo-label-assisted training (PLAT) module is adopted to reasonably refine LR labels for weakly supervised semantic segmentation of HR images. Experiments on two large-scale datasets demonstrate the superiority of Paraformer over other state-of-the-art methods for automatically updating HR land-cover maps from LR historical labels. footnotetext: 🖂 Corresponding author. The code and data are released at https://github.com/LiZhuoHong/Paraformer
@@ 3.3 Pseudo-Label-Assisted Training module
Figure 6: Six typical areas with finer observation scale on the Chesapeake Bay dataset. First row shows the LR labels (Guide). Second row shows the HR images (Source). Third row shows the HR results (Target) produced by Paraformer.
@@ 4.1 Study areas and using data
The Chesapeake Bay dataset is sampled from the largest estuary in the USA, covering 160,000 $km^{2}$ [37] and organized into 732 non-overlapping tiles, where each tile has a size of 6000 × 7500 pixels. The specific data includes:
@@
    The HR ground truths are from the OpenEarthMap [51] dataset with five land-cover classes.
Figure 8: Example of training data and different outputs of Paraformer sampled from the Chesapeake Bay dataset with 4 unified classes. (a) HR images. (b) LR labels. (c) the primal prediction from the CNN branch. (d) Mask label, as the intersection parts of (b) and (c). The black areas are set to void without supervised information. (e–f) the incorrect samples (with pink color) of LR label and mask label. (g) the final results of Paraformer. (h) HR ground truth.
@@ 4.2 Implementation Detail and Metrics
In the experiments, all methods only take LR land-cover data as training labels. Paraformer is trained by the AdamW optimizer with a patch size of 224$\times$224 and batch size of 8. The learning rate is set to 0.01 and would decrease by 10% when the loss stopped dropping over eight epochs. The metric of mean intersection over union (mIoU) is calculated between the results and the HR ground truths after their land-cover classes are unified. The compared methods include: Random Forest (RF) is a pixel-to-pixel method widely used in large-scale land-cover mapping [7]. TransUNet [10], ConViT [18], CoAtNet [16], MobileViT [34], and EfficientViT [5] are CNN-Transformer hybrid methods for semantic segmentation. UNetformer [48] and DC-Swin [47] are dedicated CNN-Transformer methods for remote-sensing images. UNet [38], HRNet [45], and LinkNet [8] are typical CNN-based semantic segmentation methods which are widely adopted in HR land-cover mapping [37, 53, 52]. SkipFCN [26] and SSDA [43] are shallow CNN-based methods for updating 1-m land-cover change maps from 30-m labels, which won first and second place in the 2021 IEEE GRSS DFC [27]. L2HNet is a state-of-the-art method designed for weakly supervised land-cover mapping [28].
@@ 4.3 Comparison Results
Comparison on the Chesapeake Bay dataset: Table 1 and Figure 5 show the comparisons on the Chesapeake Bay dataset. From the quantitative results, Paraformer shows superiority in the states of Delaware, New York, Maryland, and Pennsylvania. The L2HNet shows better results in Virginia and West Virginia. On average, Paraformer has the most accurate HR land-cover mapping results over the entire area, with a mIoU of 64.65%. As shown in Figure 5 (c), the visual result of Paraformer is more consistent with the HR image compared with other methods. Unlike the fully supervised semantic segmentation task, the unmatched training pairs can cause serious misguidedness during the model training. E.g., as the rough results shown in Figure 5 (f) and (g), UNet and HRNet over-downsample the features and encourage results to fit LR labels instead of being consistent with the HR images. Furthermore, quantitative results reveal that UNet, LinkNet, and HRNet have insufficient performance, with mIoU of 54.33%, 53.84%, and 52.00%. Although the compared CNN-Transformer methods (e.g., TransUNet) combine local and global contextual information, the structure does not focus on preserving the feature resolution or dealing with the geospatial mismatch. As a result, TransUNet shows a weak performance in visual results, shown in Figure 5 (h), and has a mIoU of 56.49%. Furthermore, SkipFCN, SSDA, and RF use small receptive fields or pixel-to-pixel strategies to extract features with fine land details. However, due to the lack of deep-level feature representation and global contextual information, SkipFCN, SSDA, and RF obtain a mIoU of 59.99%, 55.15%, and 54.56%, respectively. As an example shown in Figure 5 (e), RF finely predicts ground details but incorrectly classifies rivers, lakes, and pastures.
Comparison on the Poland dataset: In the experiments with the Poland dataset, all methods were used to produce 0.25/0.5-m land-cover maps of 14 provinces of Poland by exploiting four LR labels separately. These LR labels include 10-m FROM GLC10, ESA GLC10, Esri GLC10, and 30-m GLC FCS30. As shown in Table 2, Paraformer is compared with eight representative methods (i.e., weakly supervised, CNN-Transformer, CNN-based, pixel-to-pixel approaches) in a more extreme geospatial mismatch. Compared with the state-of-the-art method, the Paraformer has an increase in mIoU of 6.42%, 3.06%, and 4.29% in exploiting 10-m labels. By resolving 30-m labels with a max resolution gap of 120 $\times$, Paraformer has a mIoU of 49.39% with an increase of 5.77% compared with L2HNet. The typical CNN-based method has an average mIoU of 46.71% among the 10-m cases and 41.46% in the 30-m case. Skip FCN and RF have the lowest mIoU among all methods, which shows the difficulty of dealing with extremely unmatched situations. Moreover, the quantitative results of Paraformer shown in the four cases reveal that the proposed framework obtains stable results from different LR labels.
@@
Figure 7 shows the visual results of Paraformer among four cases. With the parallel CNN-Transformer structure and PLAT module, Paraformer is able to refine the clear ground details (e.g., vegetation and roads) even if they are roughly labeled in local areas. In general, Paraformer shows the potential to robustly update large-scale HR land-cover maps from available LR historical labels.
@@ 4.4 Ablation experiments
By ablating the PLAT module, the results obtained an average mIoU of 62.81%, which indicates a 1.84% decrease compared with the 64.65% of Paraformer. By ablating the CNN and Transformer branches, the results of the sole CNN branch obtained a mIoU of 60.15% and had a 4.5% decrease. Results of the sole Transformer branch obtained the lowest mIoU of 56.49% and had the most obvious decrease (8.16%). Figure 8 shows different outputs of Paraformer, where the inexact LR labels are gradually refined during framework training. The final result shown in Figure 8 (g) indicates both fine ground details and accurate land-cover patterns that are consistent with the ground truth. Moreover, Figure 9 shows the visualized contexts captured by the CNN branch, Transformer branch, and hybrid structure. Figure 9 (b) indicates that the CNN branch mostly focuses on capturing local details (e.g., the edges of roads, single houses, and shrubs). Figure 9 (c) indicates that the Transformer branch captures the feature in object scale, focusing on intact land objects of building areas and parking spots. The hybrid structure shows a strong response to the obvious objects with both fine edges and intact areas. In general, the ablation results demonstrate two findings: (1) The PLAT module can stably optimize the framework training and reasonably exploit the weak LR labels during the large-scale HR land-cover mapping process. (2) The parallel CNN and Transformer branches are indispensable parts of the framework, which construct a more robust feature extractor to bridge local and global contextual information.
@@ 5 Conclusion
In this paper, a weakly supervised CNN-Transformer framework, Paraformer, is proposed to update large-scale HR land-cover maps in an HR-label-free, end-to-end manner. Experiments on two datasets show that Paraformer outperforms other approaches in guiding semantic segmentation of large-scale HR remote-sensing images with easy-access LR land-cover data. Further analysis reveals that Paraformer can robustly adapt various landforms of wide-span areas and stably exploit different LR labels in producing accurate HR land-cover maps. The ablation studies demonstrate the effectiveness of the parallel CNN-Transformer structure and the PLAT module. Moreover, intermediate results of each training process and visualized contexts of each branch are demonstrated to transparently explain the components of Paraformer. In general, the proposed Paraformer has the potential to become an effective method for facilitating large-scale HR land-cover mapping.

### 23 Mar 2024 v3

# cvprbluergb0.21,0.49,0.74

@@ Abstract
Large-scale high-resolution (HR) land-cover mapping is a vital task to survey the Earth’s surface and resolve many challenges facing humanity. However, it is still a non-trivial task hindered by complex ground details, various landforms, and the scarcity of accurate training labels over a wide-span geographic area. In this paper, we propose an efficient, weakly supervised framework (Paraformer) to guide large-scale HR land-cover mapping with easy-access historical land-cover data of low resolution (LR). Specifically, existing land-cover mapping approaches reveal the dominance of CNNs in preserving local ground details but still suffer from insufficient global modeling in various landforms. Therefore, we design a parallel CNN-Transformer feature extractor in Paraformer, consisting of a downsampling-free CNN branch and a Transformer branch, to jointly capture local and global contextual information. Besides, facing the spatial mismatch of training data, a pseudo-label-assisted training (PLAT) module is adopted to reasonably refine LR labels for weakly supervised semantic segmentation of HR images. Experiments on two large-scale datasets demonstrate the superiority of Paraformer over other state-of-the-art methods for automatically updating HR land-cover maps from LR historical labels. footnotetext: $*{}^{*}$ndicates equal contribution. ${}^{\dagger}$orresponding author. The code and data are released at https://github.com/LiZhuoHong/Paraformer
@@ 3.3 Pseudo-Label-Assisted Training module
Figure 6: Six typical areas with finer observation scale on the Chesapeake Bay dataset. The first row shows the LR labels (Guide). The second row shows the HR images (Source). Third row shows the HR results (Target) produced by Paraformer.
@@ 4.1 Study areas and using data
The Chesapeake Bay dataset is sampled from the largest estuary in the USA and organized into 732 non-overlapping tiles, where each tile has a size of 6000 × 7500 pixels [37]. The specific data includes:
@@
    The HR ground truths are from the OpenEarthMap [51] dataset with seven land-cover classes.
Figure 8: Example of training data and different outputs of Paraformer sampled from the Chesapeake Bay dataset with four unified classes. (a) HR images. (b) LR labels. (c) the primal prediction from the CNN branch. (d) Mask label, as the intersection parts of (b) and (c). The black areas are set to void without supervised information. (e–f) the incorrect samples (with pink color) of LR label and mask label. (g) the final results of Paraformer. (h) HR ground truth.
@@ 4.2 Implementation Detail and Metrics
In the experiments, all methods only take LR land-cover data as training labels. Paraformer is trained by the AdamW optimizer with a patch size of 224$\times$224 and batch size of 8. The learning rate is set to 0.01 and would decrease by 10% when the loss stopped dropping over eight epochs. The metric of mean intersection over union (mIoU) is calculated between the results and the HR ground truths after their land-cover classes are unified into four base classes. The compared methods include: Random Forest (RF) is a pixel-to-pixel method widely used in large-scale land-cover mapping [7]. TransUNet [10], ConViT [18], CoAtNet [16], MobileViT [34], and EfficientViT [5] are CNN-Transformer hybrid methods for semantic segmentation. UNetformer [48] and DC-Swin [47] are dedicated CNN-Transformer methods for remote-sensing images. UNet [38], HRNet [45], and LinkNet [8] are typical CNN-based semantic segmentation methods which are widely adopted in HR land-cover mapping [37, 53, 52]. SkipFCN [26] and SSDA [43] are shallow CNN-based methods for updating 1-m land-cover change maps from 30-m labels, which won first and second place in the 2021 IEEE GRSS DFC [27]. L2HNet is a state-of-the-art method designed for weakly supervised land-cover mapping [28].
@@ 4.3 Comparison Results
Comparison on the Chesapeake Bay dataset: Table 1 and Figure 5 show the comparisons on the Chesapeake Bay dataset. From the quantitative results, Paraformer shows superiority in the states of Delaware, New York, Maryland, and Pennsylvania. The L2HNet shows better results in Virginia and West Virginia. On average, Paraformer has the most accurate HR land-cover mapping results over the entire area, with a mIoU of 64.65%. As shown in Figure 5 (c), the visual result of Paraformer is more consistent with the HR image compared with other methods. Unlike the fully supervised semantic segmentation task, the unmatched training pairs can cause serious misguidedness during the model training. E.g., as the rough results shown in Figure 5 (f) and (g), UNet and HRNet over-downsample the features and encourage results to fit LR labels instead of being consistent with the HR images. Furthermore, quantitative results reveal that UNet, LinkNet, and HRNet have insufficient performance, with mIoU of 54.34%, 53.84%, and 52.00%. Although the compared CNN-Transformer methods (e.g., TransUNet) combine local and global contextual information, the structure does not focus on preserving the feature resolution or dealing with the geospatial mismatch. As a result, TransUNet shows a weak performance in visual results, shown in Figure 5 (h), and has a mIoU of 56.49%. Furthermore, SkipFCN, SSDA, and RF use small receptive fields or pixel-to-pixel strategies to extract features with fine land details. However, due to the lack of deep-level feature representation and global contextual information, SkipFCN, SSDA, and RF obtain a mIoU of 59.99%, 55.15%, and 54.56%, respectively. As an example shown in Figure 5 (e), RF finely predicts ground details but incorrectly classifies rivers, lakes, and pastures.
@@
Comparison on the Poland dataset: In the experiments with the Poland dataset, all methods were used to produce 0.25/0.5-m land-cover maps of 14 provinces of Poland by exploiting four LR labels separately. These LR labels include 10-m FROM GLC10, ESA GLC10, Esri GLC10, and 30-m GLC FCS30. As shown in Table 2, Paraformer is compared with eight representative methods (i.e., weakly supervised, CNN-Transformer, CNN-based, pixel-to-pixel approaches) in a more extreme geospatial mismatch. Compared with the state-of-the-art method, the Paraformer has an increase in mIoU of 6.42%, 3.06%, and 4.29% in exploiting 10-m labels. By resolving 30-m labels with a max resolution gap of 120 $\times$, Paraformer has a mIoU of 49.39% with an increase of 5.77% compared with L2HNet. The typical CNN-based method has an average mIoU of 46.71% among the 10-m cases and 41.46% in the 30-m case. Skip FCN and RF have the lowest mIoU among all methods, which shows the difficulty of dealing with extremely unmatched situations. Moreover, the quantitative results of Paraformer shown in the four cases reveal that the proposed framework obtains stable results from different LR labels. Figure 7 shows the visual results of Paraformer among four cases. With the parallel CNN-Transformer structure and PLAT module, Paraformer is able to refine the clear ground details (e.g., vegetation and roads) even if they are roughly labeled in local areas. In general, Paraformer shows the potential to robustly update large-scale HR land-cover maps from available LR historical labels.
@@ 4.4 Ablation experiments
By ablating the PLAT module, the results obtained an average mIoU of 62.81%, which indicates a 1.84% decrease compared with the 64.65% of Paraformer. By ablating the CNN and Transformer branches, the results of the sole CNN branch obtained a mIoU of 60.15% and had a 4.5% decrease. Results of the sole Transformer branch obtained the lowest mIoU of 56.49% and had the most obvious decrease (8.16%). Figure 8 shows different outputs of Paraformer, where the inexact LR labels are gradually refined during framework training. The final result shown in Figure 8 (g) indicates both fine ground details and accurate land-cover patterns that are consistent with the ground truth. Moreover, Figure 9 shows the visualized contexts captured by the CNN branch, Transformer branch, and hybrid structure. Figure 9 (b) indicates that the CNN branch mostly focuses on capturing local details (e.g., the edges of roads, single houses, and shrubs). Figure 9 (c) indicates that the Transformer branch captures the feature in object scale, focusing on intact land objects of building areas and parking spots. The hybrid structure shows a strong response to the obvious objects with both fine edges and intact areas.
In general, the ablation results demonstrate two findings: (1) The PLAT module can stably optimize the framework training and reasonably exploit the LR labels during the large-scale HR land-cover mapping process. (2) The parallel CNN and Transformer branches are indispensable parts of the framework, which construct a more robust feature extractor to bridge local and global contextual information.
@@ 5 Conclusion
In this paper, a weakly supervised CNN-Transformer framework, Paraformer, is proposed to update large-scale HR land-cover maps in an HR-label-free, end-to-end manner. Experiments on two datasets show that Paraformer outperforms other approaches in guiding semantic segmentation of large-scale HR remote-sensing images with easy-access LR land-cover data. Further analysis reveals that the Paraformer can robustly adapt various landforms of wide-span areas and stably exploit different LR labels in producing accurate HR land-cover maps. The ablation studies demonstrate the effectiveness of the parallel CNN-Transformer structure and the PLAT module. Moreover, intermediate results of each training process and visualized contexts of each branch are demonstrated to transparently explain the components of Paraformer. In general, the proposed Paraformer has the potential to become an effective method for facilitating large-scale HR land-cover mapping.